OpenStudy (anonymous):

Write the point-slope form of an equation of the line that passes through (4, -3) and (2, 1).

OpenStudy (anonymous):

y=2x-8+-3 y=2x-11 2x-y-11=0

OpenStudy (anonymous):

A. y + 3 = - 2(x - 4) B y - 4 = - 2(x + 3) C y - 3 = - 2(x - 4)

OpenStudy (anonymous):

Do you know what the slope of your graph is?

OpenStudy (anonymous):

heres option D

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Have @satellite73 help you. I'm not sure exactly what your question is asking.

OpenStudy (anonymous):

ok thanks tho

OpenStudy (anonymous):

No problem. Good luck!

OpenStudy (anonymous):

@satellite73 help please?

OpenStudy (anonymous):

or @nincompoop